OpenStudy (anonymous):

Sean and Kyle have 26 dollars between them. If Sean has 8 more dollars than Kyle, how much money does Kyle have, in dollars? A. 18 B. –4 C. 20 D. 9

OpenStudy (anonymous):

First, set up an equation - s (sean) and k (kyle) k + s = 26 s = k+8 since we found the value of S, we can now substitute it in. k + k + 8 = 26 2k+8 = 26. Now, subtract 8 from both sides. 2k = 16 (26-8 = 16) Now, divide 16 by 2. 16/2 = 8. Kyle has 8 dollars, and Sean has 16.

OpenStudy (anonymous):

so would it be a, b, c. or d?